OpenStudy (anonymous):

find mid point.i need the equation for (8,1),(2,3) stuck on this one

OpenStudy (amistre64):

8-2 = 6 6/2 = 3 2+3 = 5 2 ia midway between 1 and 3 3,5

OpenStudy (amistre64):

typoed

OpenStudy (amistre64):

5,2 lol

OpenStudy (amistre64):

there are 6 spots between 8 and 2 right? so halfway must be 3 spots. 2 + 3 = 5; 8 - 3 = 5 right?

OpenStudy (anonymous):

k had the answer but confused on the substitute part thanks for breaking it down for me.

OpenStudy (amistre64):

the easy way for some large distances is to add the two numbers together and divide by 2 8,1 2,3 ---- 10,4 (/2) = 5,2

OpenStudy (anonymous):

thanks